OpenStudy (anonymous):

add: -5+ (-7) + (+6)

OpenStudy (king):

-6

OpenStudy (tommo_lcfc):

This is -5 -7 +6 = -12 + 6 = -6

OpenStudy (king):

-5+(-7)=-12+6=-6